QUESTION 1. Suppose Zambia is a closed economy in the Long run described by the following equations:

Y= C+I+G 

Y=5000

C= 200+0.75 (Y-T)

I= 750- 25r 

T= 1000

G= 800

Where Y is national output/income, C is consumption, I is investment, r is interest rate, G is government spending and T is taxes. 

(a) Solve for private saving, public saving and national saving.(2 marks)

(b) Solve for the equilibrium interest rate, and the level of investment (2 marks)

(c) Plot the equilibrium condition: Savings=investment (1 marks)

(d) Now suppose taxes are increased to 875:

(i) Calculate the new level of public savings, private savings, national savings, interest rate and investment.(2 marks)

(ii) Explain in few words the effect of this policy on savings, investment and interest rate. (2 marks)

(ii)Graph the new equilibrium (together with the old) for the economy due to the fiscal policy (1 marks)



1
Expert's answer
2021-02-16T11:27:28-0500

"y=C+I+G"

"C=200+0.75(5000-1000)"

"C=3200"

National saving= "Y-C-G"

"5000-3200-800=1000"

Public saving="T-G"

"1000-800=200"

Private saving="Y-T-C"

"5000-1000-3200=800"

b .Equilibrium interest (r)

"5000=3200+750-25r+800"

"250=-25r"

"r=-10"

level of investment

"750+250=1000"


c. saving=investments

1000=1000

800=1000

200=1000

d."Y=C+I+G"

"200+0.75(5000-875)=3294"

National saving="906"

public saving="75"

private saving= "831"

"5000=3294+750-25r+800"

r=-6

investments

"750-25.-6=900"


ii. Saving and investment is a function of interest